# Non-degenerate mixed maps and contact structures

### v1

@@ 3. Contact structures and open books
3.1. Introduction
Let $M$ be a closed orientable odd-dimensional manifold. A contact structure on $M$ is a field $\xi$ of hyperplanes given locally as the kernel $\xi=\operatorname{Ker}(\alpha)$ of a $1$-form $\alpha$ satisfying $\alpha\wedge(d\alpha)^{n}\neq 0$. In other words, $\xi$ is a maximally non-integrable distribution of codimension $1$. The form $\alpha$ is called a contact form. We denote $M$ endowed with this structure by $(M,\xi)$. Moreover, each contact form $\alpha$ is associated with the so-called Reeb vector field $R_{\alpha}$, uniquely determined by the following equations:
@@ 3.1. Introduction
We state the following convention. On the spheres $\mathbb{S}_{r}^{2n-1}$ and the links admitting a contact structure, the positive orientation is that given by the volume form $\lambda$ in (3.1).
Recall that two contact manifolds $(M_{1},\xi_{1})$ and $(M_{2},\xi_{2})$ are contactomorphic, or isomorphic, if there exists a diffeomorphism $\phi:M_{1}\longrightarrow M_{2}$ such that $d\phi(\xi_{1})=\xi_{2}$. Varchenko in [37] showed that the isotopy type of the contact manifold $K_{V}$ constructed from a complex isolated singularity germ $V$ does not depend on the embedding and the radius $r$ of the sphere given by the strictly plurisubharmonic function. Henceforth we shall denote the link of a variety $V$ or map germ $F$ with an isolated singularity at the origin by $K_{V}$ or $K_{F}$, respectively. An oriented contact manifold contactomorphic to such a holomorphic link is called Milnor fillable. This name is a reference to the fact that a complex link endowed with the natural contact structure is the boundary of the Milnor fiber with its natural symplectic structure. For details, see [30, Section 6].
@@
For instance, on a 3-dimensional closed oriented manifold, any contact structure is carried by some open book. Moreover, two positive contact structures carried by the same open book are isotopic. Additionally, any Milnor fillable oriented 3-manifold admits a unique Milnor fillable contact structure up to contactomorphism. Also, in dimension 3, contact structures are divided into two types: overtwisted and tight. For instance, Milnor fillable manifolds and spheres endowed with the natural structure are the first examples of tight structures. For details, see [30, Theorem 5.21 and Section 6]. A classification of overtwisted structures is developed in [10].
Let $(V,0)\subset\mathbb{C}^{n}$ be a complex germ set with an isolated singularity at the origin. Let $K_{V}=V\cap\mathbb{S}_{r}^{2n-1}$ be the link, where $r>0$ is sufficiently small. For any holomorphic function germ $h:(V,0)\longrightarrow(\mathbb{C},0)$ with an isolated singularity at the origin, one can consider the argument function $\Theta_{h}:=h/\|h\|:K_{V}\setminus h^{-1}(0)\longrightarrow\mathbb{S}^{1}$. The authors showed in [6, Theorem 3.9] that it is adapted to the natural contact structure on $K_{V}$. We remark that the underlying constructions in this theorem depend on the holomorphic setting and one cannot expect an analogous statement for the real case (see [6, Lemmas 3.6 and 3.7]). Therefore, additional hypotheses are needed to obtain open books on mixed ICIS (see Subsection 3.3).
@@ 3.2. Links of mixed ICIS
    $d\rho\wedge\alpha={\bf i}\sum_{i,j}A_{i,j}dz_{i}\wedge d{}\mkern 3.0mu\overline{\mkern-3.0muz}_{j},$
    where $A_{i,j}=2{}\mkern 3.0mu\overline{\mkern-3.0muz}_{i}z_{j}$.
@@
    $dg\wedge dh={\bf i}\sum_{i,j}B_{i,j}dz_{i}\wedge d{}\mkern 3.0mu\overline{\mkern-3.0muz}_{j}+R,$
@@
    $B_{i,j}=\frac{1}{2}\left(f_{z_{i}}\overline{f_{z_{j}}}-\overline{f_{{}\mkern 3.0mu\overline{\mkern-3.0muz}_{i}}}f_{{}\mkern 3.0mu\overline{\mkern-3.0muz}_{j}}\right).$
@@
    $d\rho\wedge\alpha\wedge dg\wedge dh=-\sum_{i,j}C_{i,j}dz_{i}\wedge d{}\mkern 3.0mu\overline{\mkern-3.0muz}_{i}\wedge dz_{j}\wedge d{}\mkern 3.0mu\overline{\mkern-3.0muz}_{j}+S,$
@@
    $C_{i,j}=\|{}\mkern 3.0mu\overline{\mkern-3.0muz}_{i}f_{z_{j}}-{}\mkern 3.0mu\overline{\mkern-3.0muz}_{b}f_{z_{i}}\|^{2}-\|z_{i}f_{{}\mkern 3.0mu\overline{\mkern-3.0muz}_{i}}-z_{j}f_{{}\mkern 3.0mu\overline{\mkern-3.0muz}_{i}}\|^{2}.$
@@
    $d\rho\wedge\alpha\wedge d\alpha^{n-2}\wedge dg\wedge dh(z,{}\mkern 3.0mu\overline{\mkern-3.0muz})$ $=\kappa(n)C(z,{}\mkern 3.0mu\overline{\mkern-3.0muz})dz_{1}\wedge d{}\mkern 3.0mu\overline{\mkern-3.0muz}_{1}\wedge\dots\wedge dz_{n}\wedge d{}\mkern 3.0mu\overline{\mkern-3.0muz}_{n},$
    where $\kappa(n)={\bf i}^{n}2^{n-2}(n-2)!$ and $C(z,{}\mkern 3.0mu\overline{\mkern-3.0muz})=\sum_{1\leq i<j\leq n}C_{i,j}$.
@@
Let $f^{1},\dots,f^{k}:(\mathbb{C}^{n},0)\longrightarrow(\mathbb{C},0)$ be mixed function germs. The following equality verifies:
$d\rho\wedge\alpha\wedge df_{1}^{1}\wedge df_{2}^{1}\wedge\dots\wedge df_{1}^{k}\wedge df_{2}^{k}=\sum_{j_{1},\dots,j_{k}}D_{j_{1},\dots,j_{k}}dz_{j_{1}}\wedge d{}\mkern 3.0mu\overline{\mkern-3.0muz}_{j_{1}}\wedge\dots\wedge dz_{j_{k}}\wedge d{}\mkern 3.0mu\overline{\mkern-3.0muz}_{j_{k}}+T,$
where $T$ is a linear combination of other types of $2k+2$-forms and
$D_{j_{1},\dots,j_{k}}$ $=2^{-k+2}\cdot C_{j_{1},j_{2}}B^{1}_{j_{3},j_{3}}\dots B^{k-1}_{j_{k},j_{k}}$
$=\left(\|{}\mkern 3.0mu\overline{\mkern-3.0muz}_{j_{1}}f^{1}_{z_{j_{2}}}-{}\mkern 3.0mu\overline{\mkern-3.0muz}_{j_{2}}f^{1}_{z_{j_{1}}}\|^{2}-\|z_{j_{1}}f^{1}_{{}\mkern 3.0mu\overline{\mkern-3.0muz}_{j_{2}}}-z_{j_{2}}f^{1}_{{}\mkern 3.0mu\overline{\mkern-3.0muz}_{j_{1}}}\|^{2}\right)\left(\|f^{2}_{z_{j_{3}}}\|^{2}-\|f^{2}_{{}\mkern 3.0mu\overline{\mkern-3.0muz}_{j_{3}}}\|^{2}\right)\dots\left(\|f^{k-1}_{z_{j_{k}}}\|^{2}-\|f^{k-1}_{z_{j_{k}}}\|^{2}\right).$
@@
We operate the wedge product of the form $d\rho\wedge\alpha\wedge df_{1}^{1}\wedge df_{2}^{1}$ with the forms $df_{1}^{m}\wedge df_{2}^{m}$ putting together the terms $dz_{i}\wedge d{}\mkern 3.0mu\overline{\mkern-3.0muz}_{i}$. Moreover, the form $T$ is the result of the wedge product of $S$ and terms with the form $B_{i,j}^{m}dz_{i}\wedge d{}\mkern 3.0mu\overline{\mkern-3.0muz}_{j}$, where $B_{i,j}^{m}$ is the expression associated to the coordinate function $f^{m}$. ∎
Let us denote the sum of $D_{j_{1},\dots,j_{k}}$ with $j_{1},\dots,j_{k}$ distinct by $D(z,{}\mkern 3.0mu\overline{\mkern-3.0muz})$. Notice that if $k=1$, it is nothing but the sum $C(z,{}\mkern 3.0mu\overline{\mkern-3.0muz})$ in item 4 of Lemma 3.4.
@@
$d\rho\wedge\alpha\wedge d\alpha^{n-(k+1)}\wedge df_{1}^{1}\wedge df_{2}^{1}\wedge\dots\wedge df_{1}^{k}\wedge df_{2}^{k}=\kappa(n)D(z,{}\mkern 3.0mu\overline{\mkern-3.0muz})dz_{1}\wedge d{}\mkern 3.0mu\overline{\mkern-3.0muz}_{1}\wedge\dots\wedge dz_{n}\wedge d{}\mkern 3.0mu\overline{\mkern-3.0muz}_{n},$
where $\kappa(n)$ is a constant depending only on $n$.
@@
We have that $d\alpha=-2i\sum_{s=1}^{n}dz_{s}\wedge d{}\mkern 3.0mu\overline{\mkern-3.0muz}_{s}$ and so
$(d\alpha)^{n-(k+1)}=\sum_{s_{1},\dots,s_{n-(k+1)}}dz_{1}\wedge d{}\mkern 3.0mu\overline{\mkern-3.0muz}_{1}\wedge\dots\wedge d\widehat{z}_{s_{1}}\wedge d\widehat{{}\mkern 3.0mu\overline{\mkern-3.0muz}}_{s_{1}}\wedge\dots\wedge d\widehat{z}_{s_{n-(k+1)}}d\widehat{{}\mkern 3.0mu\overline{\mkern-3.0muz}}_{s_{n-(k+1)}}\wedge\dots\wedge dz_{n}\wedge d{}\mkern 3.0mu\overline{\mkern-3.0muz}_{n},$
where $\widehat{z}_{s_{j}}$ are removed variables. Notice that the permutation of a pair $dz_{s_{j}}\wedge d{}\mkern 3.0mu\overline{\mkern-3.0muz}_{s_{j}}$ does not change the sign of the form. Moreover, there are (n(k+1))!(n-(k+1))! terms in the sum above and we claim that $(d\alpha)^{n-(k+1)}\wedge T=0$. Indeed, a form that appears in the sum $T$ must involve at least $k+1$ indices and, applying the wedge product with the terms of $(d\alpha)^{n-(k+1)}$, we always find repetitions and the assertion follows. ∎
Likewise in [28, Section 3.4], we have the following definition.
@@
A mixed map germ $F:(\mathbb{C}^{n},0)\longrightarrow(\mathbb{C}^{k},0)$ is called holomorphic-like (respectively anti-holomorphic-like) if $D(z,{}\mkern 3.0mu\overline{\mkern-3.0muz})\geq 0$ (respectively $D(z,{}\mkern 3.0mu\overline{\mkern-3.0muz})\leq 0$) for a sufficiently small neighborhood of the origin. If the inequalities are strict, then we call $D(z,{}\mkern 3.0mu\overline{\mkern-3.0muz})$ strictly (anti-)holomorphic-like.
Lemma 3.8.
Let $F=(f^{1},\dots,f^{k}):\mathbb{C}^{n}\longrightarrow\mathbb{C}^{2}$ be a holomorphic map germ and $\phi_{a,b}$ a homogeneous mixed covering, where $a>b$ (respectively, $a<b)$. Then the mixed map $G(w,{}\mkern 3.0mu\overline{\mkern-3.0muw})=\phi_{a,b}^{*}F$ is holomorphic-like (respectively, anti-holomorphic).
@@
Let us denote $\phi:=\phi_{a,b}$. For each $j$, the following equations hold:
$g^{l}_{w_{j}}$ $=f^{l}_{w_{j}}\left(\phi(w,{}\mkern 3.0mu\overline{\mkern-3.0muw})\right)aw_{j}^{a-1}{}\mkern 3.0mu\overline{\mkern-3.0muw}_{j}^{b},$
$g^{l}_{{}\mkern 3.0mu\overline{\mkern-3.0muw}_{j}}$ $=f^{l}_{w_{j}}\left(\phi(w,{}\mkern 3.0mu\overline{\mkern-3.0muw})\right)bw_{j}^{a}{}\mkern 3.0mu\overline{\mkern-3.0muw}_{j}^{b-1}.$
Moreover:
$\|g^{l}_{w_{j}}\|^{2}-\|g^{l}_{{}\mkern 3.0mu\overline{\mkern-3.0muw}_{j}}\|^{2}=(a^{2}-b^{2})\|\phi^{*}f^{l}_{w_{j}}\|^{2}\|w_{j}\|^{2(a+b-1)}.$
Substituting this in (3.5), the real function $D_{j_{1},\dots,j_{k}}$ is equal to
$(a^{2}-b^{2})^{k}\|w_{j_{3}}\dots w_{j_{k}}\|^{2(a+b-1)}\|w_{j_{1}}w_{j_{2}}\|^{2}\|w_{j_{1}}^{a-1}{}\mkern 3.0mu\overline{\mkern-3.0muw}_{j_{1}}^{b-1}\phi^{*}f_{z_{j_{1}}}-w_{j_{2}}^{a-1}w_{j_{2}}^{b-1}\phi^{*}f_{z_{j_{2}}}\|^{2}\|\phi^{*}f^{2}_{z_{j_{3}}}\|^{2}\dots\|\phi^{*}f^{k-1}_{z_{j_{k}}}\|^{2}.$
Theorem 3.9.
Let $F:(\mathbb{C}^{n},0)\longrightarrow(\mathbb{C}^{k},0)$ be a partially non-degenerate holomorphic map germ as in Theorem 2.10 and $\phi_{a,b}$ a homogeneous mixed covering with $a>b$ (respectively, $a<b)$. Then the link $K_{G}$ of the mixed ICIS given by $G=\phi_{a,b}^{*}F$ is a positive (respectively, negative) contact submanifold of the sphere for every sufficiently small $r>0$.
Proof.
We follow the proof of [28, Theorem 1]. We consider the case $a>b$ and the other is analogous. Theorem 2.10 implies that the link $K_{G}$ is a real smooth manifold of codimension $2k+1$ for sufficiently small $r>0$. Let us denote $G=(g^{1},\dots,g^{k})$ and $g_{1}^{i},g_{2}^{i}$ the real and imaginary parts of $g^{i}$, respectively. Notice that $K_{G}$ is a complete intersection defined by $g_{1}^{i}=g_{2}^{i}=0$ and $\rho-r^{2}=0$, for $i=1,\dots,k$. This implies that there exists a local coordinate system formed by $g_{1}^{i},g_{2}^{i},\rho$, and other real analytic functions $h_{2k+2},\dots,h_{2n}$. Therefore, the condition $\alpha\wedge(d\alpha)^{n-(k+1)}\not\equiv 0$ is equivalent to
$d\rho\wedge\alpha\wedge d\alpha^{n-(k+1)}\wedge dg_{1}^{1}\wedge dg_{2}^{1}\wedge\dots\wedge dg_{1}^{k}\wedge dg_{2}^{k}\not\equiv 0.$
[…]
$\begin{cases}&\|w_{j_{3}}\dots w_{j_{k}}\|^{2(a+b-1)}\|w_{j_{1}}w_{j_{2}}\|^{2}\|w_{j_{1}}^{a-1}{}\mkern 3.0mu\overline{\mkern-3.0muw}_{j_{1}}^{b-1}\phi^{*}f_{z_{j_{1}}}-w_{j_{2}}^{a-1}w_{j_{2}}^{b-1}\phi^{*}f_{z_{j_{2}}}\|^{2}\|\phi^{*}f^{2}_{z_{j_{3}}}\|^{2}\dots\|\phi^{*}f^{k-1}_{z_{j_{k}}}\|^{2}\equiv 0\\ &\phi^{*}f^{j}\equiv 0\;\,\text{for all}\;j=1,\dots,k,\end{cases}$
where $f^{i}$ and $f_{z_{j_{l}}}^{i}$ are restricted to the curve $w(t)$. Let $I=\{j:w_{j}(t)\not\equiv 0\}$. By the non-degeneracy condition of the coordinate functions restricted to $V_{G}$ and Lemma 2.2, there exist $j_{1},\dots,j_{k}\in I$ such that $f^{l}_{z_{j_{k}}}(w(t))\not\equiv 0$ for each $l=1,\dots,k$. Then $D_{j_{1},\dots,j_{k}}\equiv 0$ and
(3.3) $\left\|w_{j_{1}}^{a-1}{}\mkern 3.0mu\overline{\mkern-3.0muw}_{j_{1}}^{b-1}\phi^{*}f^{1}_{z_{j_{1}}}-w_{j_{2}}^{a-1}w_{j_{2}}^{b-1}\phi^{*}f^{1}_{z_{j_{2}}}\right\|^{2}\equiv 0.$
Let $J=\{j_{l}:w_{j_{l}}(t)\not\equiv 0\}$ and define the curve $\gamma(t)=w_{j_{l}}^{a-1}{}\mkern 3.0mu\overline{\mkern-3.0muw}_{j_{l}}^{b-1}\phi^{*}f_{z_{j_{l}}}(w(t),{}\mkern 3.0mu\overline{\mkern-3.0muw}(t))$. Note that $\gamma(t)\not\equiv 0$ and any $j_{l}\in J$ such that $f_{z_{j_{l}}}\left(\phi(w(t),{}\mkern 3.0mu\overline{\mkern-3.0muw}(t))\right)\not\equiv 0$ defines the same curve by the equation (3.3). Let $v_{j_{l}}(t)=w_{j_{l}}^{\prime}(t)$. We have that:
$0$ $=\frac{dg(w(t),{}\mkern 3.0mu\overline{\mkern-3.0muw}(t))}{dt}$
$=\sum_{l}^{n}g_{w_{j_{l}}}\frac{dw_{j_{l}}}{dt}+g_{{}\mkern 3.0mu\overline{\mkern-3.0muw}_{j}}\frac{d{}\mkern 3.0mu\overline{\mkern-3.0muw}_{j_{l}}}{dt}$
$=\sum_{l}^{n}f_{z_{j_{l}}}(\phi(w(t),w(t)))\left(aw_{j_{l}}^{a-1}{}\mkern 3.0mu\overline{\mkern-3.0muw}_{j_{l}}^{b}v_{j_{l}}+bw_{j_{l}}^{a}{}\mkern 3.0mu\overline{\mkern-3.0muw}_{j_{l}}^{b-1}{}\mkern 3.0mu\overline{\mkern-3.0muv}_{j_{l}}\right)$
$=\sum_{l}^{n}f_{z_{j_{l}}}(\phi)w_{j_{l}}^{a-1}{}\mkern 3.0mu\overline{\mkern-3.0muw}_{j_{l}}^{b-1}\left(a{}\mkern 3.0mu\overline{\mkern-3.0muw}_{j_{l}}v_{j}+bw_{j_{l}}{}\mkern 3.0mu\overline{\mkern-3.0muv}_{j_{l}}\right)$
$=\gamma(t)\sum_{l}^{n}\left(a{}\mkern 3.0mu\overline{\mkern-3.0muw}_{j_{l}}v_{j_{l}}+bw_{j}{}\mkern 3.0mu\overline{\mkern-3.0muv}_{j_{l}}\right).$
Notice that the sum above can be taken on $j_{l}\in J$, because if $j_{l}\notin J$ we have null terms. Since this last term is zero, summing up $\sum_{l}aw_{j_{l}}{}\mkern 3.0mu\overline{\mkern-3.0muv}_{j_{l}}$ and $\sum_{j_{l}}b{}\mkern 3.0mu\overline{\mkern-3.0muw}_{j_{l}}v_{j_{l}}$ does not change the equality. Thus
$0=\gamma(t)\left(\frac{a+b}{2}\right)\sum_{l}^{n}w_{j_{l}}{}\mkern 3.0mu\overline{\mkern-3.0muv}_{j_{l}}+{}\mkern 3.0mu\overline{\mkern-3.0muw}_{j_{l}}v_{j_{l}}.$
Taking the derivative of $\|w(t)\|^{2}=\sum_{l}^{n}w_{j_{l}}{}\mkern 3.0mu\overline{\mkern-3.0muw}_{j_{l}}$ we obtain:
$\frac{d\|w(t)\|^{2}}{dt}=\sum_{l}^{n}w_{j_{l}}{}\mkern 3.0mu\overline{\mkern-3.0muv}_{j_{l}}+{}\mkern 3.0mu\overline{\mkern-3.0muw}_{j_{l}}v_{j_{l}}=0.$
Since $\lim_{t\to 0}w(t)=0$, it implies $w(t)\equiv 0$, which is a contradiction and the result follows. ∎

### v3

@@ 3. Contact structures and open books
3.1. Basic facts
Let $M^{2n+1}$ be a closed orientable odd-dimensional manifold. A contact structure on $M$ is a field $\xi$ of hyperplanes given locally as the kernel $\xi=\operatorname{Ker}(\alpha)$ of a $1$-form $\alpha$ satisfying $\alpha\wedge(d\alpha)^{n}\neq 0$. In other words, $\xi$ is a maximally non-integrable distribution of codimension $1$. The form $\alpha$ is called a contact form. We denote $M$ endowed with this structure by $(M,\xi)$. Moreover, each contact form $\alpha$ is associated with the so-called Reeb vector field $R_{\alpha}$, uniquely determined by the following equations:
@@ 3.1. Basic facts
We state the following convention. On the spheres $\mathbb{S}_{r}^{2n-1}$ and the links admitting a contact structure, the positive orientation is that given by the volume form $\alpha\wedge(d\alpha)^{n}$.
Recall that two contact manifolds $(M_{1},\xi_{1})$ and $(M_{2},\xi_{2})$ are contactomorphic, or isomorphic, if there exists a diffeomorphism $\phi:M_{1}\longrightarrow M_{2}$ such that $d\phi(\xi_{1})=\xi_{2}$. Varchenko in [42] showed that the isotopy type of the contact manifold $K_{V}$ constructed from a complex isolated singularity germ $V$ does not depend on the embedding and the radius $r$ of the sphere given by the strictly plurisubharmonic function. Henceforth we shall denote the link of a variety $V$ or map germ $F$ with an isolated singularity at the origin by $K_{V}$ or $K_{F}$, respectively. An oriented contact manifold contactomorphic to such a holomorphic link is called Milnor fillable. This name is a reference to the fact that a complex link endowed with the natural contact structure is the boundary of the Milnor fiber with its natural symplectic structure. For details, see [33, Section 6].
@@
For instance, on a 3-dimensional closed oriented manifold, any contact structure is carried by some open book. Moreover, two positive contact structures carried by the same open book are isotopic. Additionally, any Milnor fillable oriented 3-manifold admits a unique Milnor fillable contact structure up to contactomorphism. Also, in dimension 3, contact structures are divided into two types: overtwisted and tight. For instance, Milnor fillable manifolds and spheres endowed with the natural structure are the first examples of tight structures. For details, see [33, Theorem 5.21 and Section 6]. A classification of overtwisted structures is developed in [12].
Let $(V,0)\subset\mathbb{C}^{n}$ be a complex germ set with an isolated singularity at the origin. Let $K_{V}=V\cap\mathbb{S}_{r}^{2n-1}$ be the link, where $r>0$ is sufficiently small. For any holomorphic function germ $h:(V,0)\longrightarrow(\mathbb{C},0)$ with an isolated singularity at the origin, one can consider the argument function $\Theta_{h}:=h/\|h\|:K_{V}\setminus h^{-1}(0)\longrightarrow\mathbb{S}^{1}$. The authors showed in [6, Theorem 3.9] that it is adapted to the natural contact structure on $K_{V}$. We remark that the underlying constructions in this theorem depend on the holomorphic setting, and one cannot expect an analogous statement for the real case (see [6, Lemmas 3.6 and 3.7]). Therefore, additional hypotheses are needed to obtain open books on links defined by mixed ICIS (see Subsection 3.3).
@@ 3.2. Links of mixed ICIS
    $d\rho\wedge\alpha={\bf i}\sum_{i,j}\mathcal{A}_{i,j}dz_{i}\wedge d{}\mkern 3.0mu\overline{\mkern-3.0muz}_{j},$
    where $\mathcal{A}_{i,j}=2{}\mkern 3.0mu\overline{\mkern-3.0muz}_{i}z_{j}$.
@@
    $dg\wedge dh={\bf i}\sum_{i,j}\mathcal{B}^{f}_{i,j}dz_{i}\wedge d{}\mkern 3.0mu\overline{\mkern-3.0muz}_{j}+R,$
@@
    $\mathcal{B}^{f}_{i,j}=\frac{1}{2}\left(f_{z_{i}}\overline{f_{z_{j}}}-\overline{f_{{}\mkern 3.0mu\overline{\mkern-3.0muz}_{i}}}f_{{}\mkern 3.0mu\overline{\mkern-3.0muz}_{j}}\right).$
@@
    $d\rho\wedge\alpha\wedge dg\wedge dh=-\sum_{i,j}\mathcal{C}_{i,j}^{f}dz_{i}\wedge d{}\mkern 3.0mu\overline{\mkern-3.0muz}_{i}\wedge dz_{j}\wedge d{}\mkern 3.0mu\overline{\mkern-3.0muz}_{j}+S,$
@@
    $\mathcal{C}_{i,j}^{f}=\|{}\mkern 3.0mu\overline{\mkern-3.0muz}_{i}f_{z_{j}}-{}\mkern 3.0mu\overline{\mkern-3.0muz}_{j}f_{z_{i}}\|^{2}-\|z_{i}f_{{}\mkern 3.0mu\overline{\mkern-3.0muz}_{j}}-z_{j}f_{{}\mkern 3.0mu\overline{\mkern-3.0muz}_{i}}\|^{2}.$
@@
    $d\rho\wedge\alpha\wedge d\alpha^{n-2}\wedge dg\wedge dh(z,{}\mkern 3.0mu\overline{\mkern-3.0muz})$ $=\kappa(n)\mathcal{C}^{f}(z,{}\mkern 3.0mu\overline{\mkern-3.0muz})dz_{1}\wedge d{}\mkern 3.0mu\overline{\mkern-3.0muz}_{1}\wedge\dots\wedge dz_{n}\wedge d{}\mkern 3.0mu\overline{\mkern-3.0muz}_{n},$
    where $\kappa(n)={\bf i}^{n}2^{n-2}(n-2)!$ and $\mathcal{C}^{f}(z,{}\mkern 3.0mu\overline{\mkern-3.0muz})=\sum_{1\leq i<j\leq n}\mathcal{C}_{i,j}^{f}$.
@@
Let $F=(f^{1},\dots,f^{k}):(\mathbb{C}^{n},0)\longrightarrow(\mathbb{C}^{k},0)$ be a mixed map germ. For each $j$, write $f^{j}=f^{j}_{1}+{\bf i}f^{j}_{2}$ for the real and imaginary parts of $f^{j}$. The following equality verifies:
$d\rho\wedge\alpha\wedge df_{1}^{1}\wedge df_{2}^{1}\wedge\dots\wedge df_{1}^{k}\wedge df_{2}^{k}=({\bf i})^{k+1}\sum_{J}\mathcal{D}_{J}^{F}dz_{j_{1}}\wedge d{}\mkern 3.0mu\overline{\mkern-3.0muz}_{j_{1}}\wedge\dots\wedge dz_{j_{k+1}}\wedge d{}\mkern 3.0mu\overline{\mkern-3.0muz}_{j_{k+1}}+T,$
where
$\mathcal{D}_{J}^{F}=\mathcal{A}_{j_{1},j_{2}}\mathcal{B}^{f^{1}}_{j_{3},j_{4}}\cdots\mathcal{B}^{f^{k}}_{j_{2k+1},j_{2k+2}}$
with $j_{1},\dots,j_{2k+2}\in J=\{j_{1},\dots,j_{k+1}\}$ a set of $k+1$ distinct indices and $T$ is a linear combination of other types of $2k+2$-forms.
@@
We operate the wedge product in such a way to produce the elements of the form $dz_{j_{1}}\wedge d{}\mkern 3.0mu\overline{\mkern-3.0muz}_{j_{1}}\wedge\dots\wedge dz_{j_{k+1}}\wedge d{}\mkern 3.0mu\overline{\mkern-3.0muz}_{j_{k+1}}$. There are several ways to obtain such elements, and thus $\mathcal{D}_{J}^{F}$ comprehends one possibility. Moreover, the form $T$ is the result of the wedge product of $S$ and terms with the form $\mathcal{B}_{i,j}^{f^{m}}dz_{i}\wedge d{}\mkern 3.0mu\overline{\mkern-3.0muz}_{j}$ or $\mathcal{B}_{i,j}^{f^{m}}d{}\mkern 3.0mu\overline{\mkern-3.0muz}_{i}\wedge d{}\mkern 3.0mu\overline{\mkern-3.0muz}_{j}$, where $\mathcal{B}_{i,j}^{f^{m}}$ is the expression associated to the coordinate function $f^{m}$ such that the resulting expression contains more than $k+1$ indices and hence does not generate the previous elements. ∎
Let us denote the sum of $\mathcal{D}_{J}^{F}$ with $j_{1},\dots,j_{k+1}$ distinct by $\mathcal{D}^{F}(z,{}\mkern 3.0mu\overline{\mkern-3.0muz})$. Notice that if $k=1$, it is nothing but the sum $\mathcal{C}^{f}(z,{}\mkern 3.0mu\overline{\mkern-3.0muz})$ in item 4 of Lemma 3.4.
@@
$d\rho\wedge\alpha\wedge d\alpha^{n-(k+1)}\wedge df_{1}^{1}\wedge df_{2}^{1}\wedge\dots\wedge df_{1}^{k}\wedge df_{2}^{k}=\kappa(n)\mathcal{D}^{F}(z,{}\mkern 3.0mu\overline{\mkern-3.0muz})dz_{1}\wedge d{}\mkern 3.0mu\overline{\mkern-3.0muz}_{1}\wedge\dots\wedge dz_{n}\wedge d{}\mkern 3.0mu\overline{\mkern-3.0muz}_{n},$
where $\kappa(n)=({\bf i})^{n}2^{n-k-1}\left(n-(k+1)\right)!$ is a constant depending only on $n$.
@@
We have that $d\alpha=2{\bf i}\sum_{s=1}^{n}dz_{s}\wedge d{}\mkern 3.0mu\overline{\mkern-3.0muz}_{s}$ and so
$(d\alpha)^{n-(k+1)}=(2{\bf i})^{n-(k+1)}\sum_{s_{1},\dots,s_{n-(k+1)}}dz_{1}\wedge d{}\mkern 3.0mu\overline{\mkern-3.0muz}_{1}\wedge\dots\wedge d\widehat{z}_{s_{1}}\wedge d\widehat{{}\mkern 3.0mu\overline{\mkern-3.0muz}}_{s_{1}}\wedge\dots\wedge d\widehat{z}_{s_{n-(k+1)}}d\widehat{{}\mkern 3.0mu\overline{\mkern-3.0muz}}_{s_{n-(k+1)}}\wedge\dots\wedge dz_{n}\wedge d{}\mkern 3.0mu\overline{\mkern-3.0muz}_{n},$
where $\widehat{z}_{s_{j}}$ are removed variables. Notice that the permutation of a pair $dz_{s_{j}}\wedge d{}\mkern 3.0mu\overline{\mkern-3.0muz}_{s_{j}}$ does not change the sign of the form. Moreover, there are (n(k+1))!(n-(k+1))! terms in the sum above and we claim that $(d\alpha)^{n-(k+1)}\wedge T=0$. Indeed, a form that appears in the sum $T$ must involve at least $k+2$ indices and, applying the wedge product with the terms of $(d\alpha)^{n-(k+1)}$, we always find repetitions, and the assertion follows. ∎
Likewise, in [32, Section 3.4], we have the following definition.
@@
A mixed ICIS $F:(\mathbb{C}^{n},0)\longrightarrow(\mathbb{C}^{k},0)$ is called holomorphic-like (respectively anti-holomorphic-like) if $\mathcal{D}^{F}(z,{}\mkern 3.0mu\overline{\mkern-3.0muz})\geq 0$ (respectively $\mathcal{D}^{F}(z,{}\mkern 3.0mu\overline{\mkern-3.0muz})\leq 0$) for a sufficiently small neighborhood of the origin. If the inequalities are strict, then we call $\mathcal{D}^{F}(z,{}\mkern 3.0mu\overline{\mkern-3.0muz})$ strictly (anti-)holomorphic-like.
Example 3.8.
For instance, by the constructions in [32], if $g=\phi^{*}_{\textbf{a},\textbf{b}}f$ is a mixed function given as the pullback of a holomorphic function by a homogeneous mixed covering, then $g$ is (respectively, anti-)holomorphic-like if $a>b$ (respectively, $a<b$). If we further require that $f$ is non-degenerate and convenient, then this property holds strictly. By Lemma 2.3, if $f$ is a holomorphic partially non-degenerate function, then the pullback by a homogeneous mixed covering is an isolated singularity. Following the argument of [32, Theorem 1], by Lemma 2.2, we see that the same proof applies, and Oka’s result can be extended to a larger class of mixed functions.
The next statement is a formulation of [32, Theorem 1] for the case of maps into $\mathbb{C}^{k}$ and justifies the previous nomenclature.
Theorem 3.9.
Let $F:(\mathbb{C}^{n},0)\longrightarrow(\mathbb{C}^{k},0)$ be a strictly (respectively, anti-)holomorphic-like mixed ICIS. Then the link $K_{F}$ of the variety $V_{F}$ defined by $F$ is a positive (respectively, negative) contact submanifold of the sphere for every sufficiently small $r>0$.
@@
We have that the link $K_{F}$ is a real smooth manifold of codimension $2k+1$ for sufficiently small $r>0$. Notice that $K_{F}$ is a complete intersection defined by $f_{1}^{i}=f_{2}^{i}=0$ and $\rho-r^{2}=0$, for $i=1,\dots,k$. This implies that there exists a local coordinate system formed by $f_{1}^{i},f_{2}^{i},\rho$, and other real analytic functions $h_{2k+2},\dots,h_{2n}$. Therefore, the condition $\alpha\wedge(d\alpha)^{n-(k+1)}\neq 0$ is equivalent to
(3.3) $d\rho\wedge\alpha\wedge d\alpha^{n-(k+1)}\wedge df_{1}^{1}\wedge df_{2}^{1}\wedge\dots\wedge df_{1}^{k}\wedge df_{2}^{k}\neq 0,$
and Corollary 3.6 now implies the statement. ∎